b1az3:

Solve x2 - 8x + 3 = 0 by completing the square. Which equation is used in the process? ( x - 4) 2 = 0 ( x - 4) 2 = 19 ( x - 4) 2 = 13 Use a table of function values to approximate an x-value in which the exponential function exceeds the polynomial function. f(x) = 3(3)(2x-5) h(x) = (x + 1)5 + 4 x = 5 x = 6 x = 7 x = 8

jhonyy9:

do you know how you make square complete from these x^2 -8x +3 = 0 ?

jhonyy9:

suppose -8x you get assuming -4+(-4) = -8 so what mean this ? how many need be the ,,c" term in place of 3 ?

jhonyy9:

hope you know the standard form of a quadratic Ax^2 +Bx +C = 0

jhonyy9:

B = -4+(-4) = -8 C = -4*(-4) = ?

jhonyy9:

C = 16

jhonyy9:

how you make from +3 this 16 ? how many you need adding to +3 to get +16 ?

b1az3:

that exam

jhonyy9:

i like make you understanding step by step @AZ

jhonyy9:

@Shadow

jhonyy9:

ok x^2 -8x +3 = 0 x*x -4x-4x +3 +13 -13 = 0 x*x -4x-4x +4*4 = 13 x(x-4) -4(x-4) = 13 (x-4)(x-4)=13 (x-4)^2 = 13 hope helped understandably easy

jhonyy9:

@b1az3

jhonyy9:

@dizzymustang156 opinion pls ? ty.

dizzymustang156:

amazing

jhonyy9:

hope you know the way how you complet a quadratic to be perfect square ...,